A waste disposal company is considering the replacement of one of its aging trucks. The key parameters of the three trucks under scrutiny are provided below.

Parameters

Delta

Epsilon

Zeta

1. Initial Cost

($)

250,000

375,000

450,000

2. Revenues

($)

230,000 at EOY1

increasing by 2.5% annually thereafter

195,000 at EOY1

increasing by 3,000 annually thereafter

235,000 at EOY1

decreasing by 1% annually thereafter

3. Operating

costs ($)

140,000 at EOY1

decreasing by 2,000 annually thereafter

125,000 at EOY1

decreasing by 2% annually thereafter

EOY1-EOY4 = 125,000;

EOY5-EOY8 = 135,000

EOY9-EOY12 = 170,000

EOY13-EOY16 = 190,000

4. End-of-life salvage value ($)

-20,000

7,000

-20,000

5. Useful life

(years)

4

8

16

  • EOY = End-of-year
  • Industry Standard = 4 years
  • MARR = 10%
  1. The best truck based on the simple payback method is
    a) Delta; b) Epsilon; c) Zeta.
  1. Delta’s benefit/cost (B/C) ratio (second decimal; no rounding) is
    a) 0.98; b) 1.03; c) 1.08; d) 1.10.
  1. Zeta’s benefit/cost (B/C) ratio (second decimal; no rounding) is
    a) 0.99; b) 1.02; c) 1.06; d) 1.09.
  1. The incremental B/C ratio (second decimal; no rounding) between the Delta and Epsilon trucks is

a) 0.90; b) 0.99; c) 1.03; d) 1.08.

Hint: Delta's NFW = $81,900

Epsilon's NPW = $89,000

Zeta's NFW = $709,100
Answer #1

Total Cost of Delta Trucks = Initial Cost + Operating Cost + Salvage Value (since salvage is negative )

                                               = 250,000 + (140,000 +138,000 + 136,000 + 134,000) + 20,000

                                               = 818,000

& Total Revenue from Delta Trucks= Revenue

                                                              = 230,000 +235750 + 241643.75 + 247684.84

                                                              = 955,078.59

                                              

Total Cost of Epsilon Trucks = Initial Cost + Operating Cost

                                               = 375,000 + (125,000 +122,500 + 120050 + 117649 + 115296 +112990                                +110730 +108515)

                                               = 1195230

& Total Revenue from Epsilon Trucks= Revenue + Salvage Value

                                                              = 195,000 +198000 + 201000 + 204000 +207000 +210000 + 213000 + 217000 + 220000 + 7000

                                                              = 1165,000

Total Cost of Zeta Trucks = Initial Cost + Operating Cost + Salvage Value (since salvage is negative )

                                               = 450,000 + (125,000 x4 + 135000x4 + 170000x4 + 190000x4) + 20000

                                               = 450000 + 500000 + 540000 + 680000 + 760000 + 20000

                                                = 2,950,000

& Total Revenue from Epsilon Trucks= Revenue + Salvage Value

                                                              = 235,000 +232500 + 230700 + 227000+ 224021+222147+220069+219451+217985+215784+213545+245615

                                                              = 3,650,000

Therefore as per the calculations above, The best truck is is (b) Epsilon

Delta's benefit cost is (b) 1.03%

Delta's benefit cost is (a) 0.99

Incremental ratio is (d) 1.08